# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ 4.4 DiD estimators with covariates that target the ATT(2)
Fortunately, TWFE is not the only way to bring covariates into DiD estimation. Alternative strategies start from the identification result in (4.2):
@@
One way to obtain the second term is to first estimate the inner conditional expectation, $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$. This object is just an (unknown) function that relates average outcome trends for untreated units to their covariates. The most common way to proceed, especially in cases where $X_{i}$ contains many variables or continuous ones, is to specify a working model, ${\mu}_{\omega,\Delta,D=0}(X_{i})$, with parameters that are simple to estimate. A natural and empirically friendly choice is a linear model, ${\mu}_{\omega,\Delta,D=0}(X_{i})=X_{i}^{\prime}\beta_{D=0}$ whose parameters come from a (weighted) regression of $\Delta Y_{i,t=2}$ on $X_{i}$ in the sample of untreated units. The results of this regression describe untreated outcome trends as a function of $X_{i}$. The fitted model then generates predicted values, $\widehat{\mu}_{\omega,\Delta,D=0}(X_{i})=X_{i}^{\prime}\widehat{\beta}_{D=0}$, for all units in the sample, including treated units. With these fitted values we can estimate $\mathop{}\!\mathbb{E}_{\omega}\Big{[}\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]\Big{|}D_{i}=1\Big{]}$ using the plug-in principle, i.e., by replacing $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$ with its fitted value $\widehat{\mu}_{\omega,\Delta,D=0}(X_{i})$, and replacing population (weighted) expectations with their sample analogs: $\frac{\sum_{i=1}^{n}D_{i}\leavevmode\nobreak\ \omega_{i}\leavevmode\nobreak\ \widehat{\mu}_{\omega,\Delta,D=0}(X_{i})}{\sum_{i=1}^{n}D_{i}\leavevmode\nobreak\ \omega_{i}}$.
Putting the pieces together gives the following estimator for $ATT(2)$
@@
This strategy is often referred to as the regression-adjustment (RA) or outcome regression approach to DiD, see, e.g., Heckman1997.
We apply this strategy in our application using only the baseline covariates from Table 4 and report the parameters $\widehat{\beta}_{D=0}$ of our working model in columns (1) and (3) of Table 6. In practice, mortality changes in non-expansion states are only weakly related to our baseline covariates. Multiplying the weighted coefficients in Table 6 times the weighted treatment group 2013 means in Table 4 gives a predicted change in untreated mortality rates for the average treated county of 7.2 deaths per 100,000, the second term in equation (4.5). A fit this close to the observed (weighted) change of 6.3 deaths indicates the poor fit of our outcome model. The observed weighted trend in mortality for expansion counties from Table 2 is 3.7 deaths. Together, these imply that this approach, based only on assumptions CPT, SO, NA, the choice of baseline covariates, and a linear model for $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$, yields an estimated $ATT(2014)$ of -3.5. This matches the formal RA DiD estimate we report in column 4 of Table 7 (labeled as “Regression”). Column 1, however, gives an unweighted estimate from the same procedure of -1.62.
@@
Unweighted Weighted Regression Propensity Score Regression Propensity Score (1) (2) (3) (4) OLS Logit OLS Logit Constant -20.91 -10.00∗∗∗ -4.62 -8.17∗∗∗ (69.85) (1.35) (44.98) (0.01) % Female 0.04 -0.04∗∗ -0.09 -0.19∗∗∗ (0.82) (0.02) (0.68) (0.00) % White 0.15 0.06∗∗∗ 0.20 0.04∗∗∗ (0.22) (0.00) (0.11) (0.00) % Hispanic -0.08 -0.02∗∗∗ -0.08 -0.02∗∗∗ (0.20) (0.00) (0.08) (0.00) Unemployment Rate 1.14 0.32∗∗∗ 0.88 0.68∗∗∗ (1.56) (0.03) (0.99) (0.00) Poverty Rate 0.21 0.03 -0.13 0.11∗∗∗ (0.98) (0.02) (0.53) (0.00) Median Income 0.09 0.08∗∗∗ -0.05 0.15∗∗∗ (0.52) (0.01) (0.24) (0.00) Notes: This table reports the outcome regression propensity score models that enter into the estimator from SantAnna2020 and Callaway2021. The first two columns report the results for unweighted regressions and the second two report results from weighted regression models. The regression model predicts changes in the outcome variable (mortality rates) on the 2013 covariate values for just the counties that do not expand Medicaid in 2014. The propensity score model uses data for 2013 and estimates a logit model of an expansion indicator variable on the 2013 covariate levels. Standard errors (in parentheses) are clustered at the county level.
@@
To compute standard errors, we need to account for the fact that (4.5) is a two-step estimation procedure and take into account the uncertainty associated with estimating the working model $\mu_{\omega,\Delta,D=0}(X_{i})$. This is standard, though, and most statistical software automates this process; see, e.g., SantAnna2020 and Callaway2021.
@@
  • Notes: This table reports the 2 $\times$ 2 DiD estimate comparing counties that expand Medicaid in 2014 to counties that do not expand Medicaid, adjusting for the inclusion of 2013 covariate values using the methodologies discussed in SantAnna2020 and Callaway2021. The first column reports results using regression adjustment, the second column uses inverse probability weighting based on a propensity score model using the included covariates, and the third column uses the doubly robust combination of the two approaches. Standard errors (in parentheses) are clustered at the county level.
A linear working model for $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$ is a familiar choice, but not the only one. More flexible, or even fully nonparametric, working models are possible, and the procedure is the same: estimate the model on untreated units, get its fitted values for the covariate values of the treated units, and then estimate the $ATT(2)$ using (4.5). An important consideration when choosing the working models is sample size. Large samples permit more flexible estimators that do not sacrifice too much precision. In smaller samples, a parametric linear model may be more appealing. Ultimately, the reliability of DiD RA estimators for $ATT(2)$ depends on how well $\widehat{\mu}_{\omega,\Delta,D=0}(X_{i})$ approximates $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$. If the working model is misspecified, for example by omitting relevant nonlinear terms, the resulting DiD RA estimator will be biased.
@@
To implement the IPW DiD procedure and construct these “balancing weights”, we need to model $p_{\omega}(X_{i})=P_{\omega}(D_{i}=1|X_{i})$, the (weighted) conditional probability of belonging to the treatment group. IPW weights are a function of $p_{\omega}(X_{i})$ and, for estimating $ATT(2)$, they take a form that forces the underlying weighted distribution of covariates for comparison units to match the distribution for treatment units (Rosenbaum1983). [^14] To learn about $ATT(2)$, we do not need to re-weight the treated units because the identification challenge is to estimate an untreated counterfactual. In fact, the RA estimator we discussed has a similar feature: it requires an outcome model for untreated outcomes only, not treated ones. Intuitively, these weights are formed such that if we find some units that were likely to be observed in the treatment groups (based on their covariate values) but ended up in the comparison group, we give these untreated observations “extra” weight.
@@
$ATT(2)$ $=\mathop{}\!\mathbb{E}\left[\Big{(}w_{\omega,D=1}(D_{i})-w_{\omega,D=0}(D_{i},X_{i})\Big{)}\Delta Y_{i,t=2}\right]$ (4.6)
@@
$w_{\omega,D=1}(D)=D\omega\Big{/}{\mathop{}\!\mathbb{E}[\omega D]},\leavevmode\nobreak\ \text{and }\leavevmode\nobreak\ w_{\omega,D=0}(D,X)=\dfrac{\omega(1-D)p_{\omega}(X)}{1-p_{\omega}(X)}\Bigg{/}\mathop{}\!\mathbb{E}\left[\dfrac{\omega(1-D)p_{\omega}(X)}{1-p_{\omega}(X)}\right],$ (4.7)
see, e.g., Abadie2005 and SantAnna2020. The structure of the weights in equation (4.7) and the way they enter equation (4.6) highlights several intuitive features of how the IPW estimator works. First, $w_{\omega,D=1}(D)$ is only non-zero for treated units, and $w_{\omega,D=0}(D,X)$ is only non-zero for untreated units, which means the estimator subtracts a particular mean of outcome trends for untreated units from a particular mean of outcome trends for treated units. Second, the $w_{\omega,D=1}(D)$ weights do not involve $X_{i}$ and simply lead simply to a ($\omega$-weighted) mean for the treatment group. Third, the $w_{\omega,D=0}(D,X)$ weights are functions of $X_{i}$ that, as (4.7) shows, give increasingly more weight to untreated units with high propensity scores. This specific weighting function is what allows IPW weights to build a comparison group whose covariate distribution matches the treatment group. Fourth, equation (4.7) clarifies that two types of weights both enter an IPW analysis. We already discussed how the $\omega$ weights shape the parameter of interest, parallel trends assumptions and estimation. These simply multiply the IPW weights which act to address imbalance (and their product is rescaled to integrate to one within group). Finally, the appeal of (4.6) is that if we have a better sense of how units sort into treatment than of the factors that shape outcome trends, we may be more comfortable modeling $p_{\omega}(X_{i})$ than $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$.
@@
Based on these fitted values, we can estimate $ATT(2)$ by
@@
We report the $ATT(2014)$ estimates and their standard errors using this IPW DiD procedure in Table 5 (labeled as “IPW”) using both unweighted and population-weighted procedures, where we use a logistic regression that is linear in covariates as our propensity score working model. We use the delta method to account for the estimation uncertainty inherited in this two-step estimation procedure when computing standard errors. The IPW DiD estimates are similar to the RA DiD estimates in that we obtain positive estimates when we do not use population weights and negative estimates when we weight. However, the weighted IPW estimate is almost twice as large as the RA estimate, despite neither being statistically significant, and the unweighted IPW estimate is less than half the size of the RA estimate. This is consistent with the broad conclusion from Table 6 that our covariates explain Medicaid expansion better than mortality trends.
IPW estimators for $ATT(2)$ tend to be noisy when fitted propensity score estimates are too close to 1 among untreated units, a condition related to the Assumption SO. The reason for this is simple: when $\widehat{\pi}_{\omega}(X_{i})$ is close to one among units with $D_{i}=0$, the (estimated) inverse probability weights $\widehat{w}_{\omega,D=0}(D,X)$ becomes more volatile, as one is essentially diving by zero. A good practice when using IPW estimators is to check the plausibility of strong overlap using estimated propensity scores. Figure 1 plots the propensity scores that come from the logit models in Table 6 for the two groups of counties. Few untreated units have very high estimated propensity scores, so extreme weighting is not a significant concern. In addition, propensity scores of non-expansion counties seem to lie within the support of the expansion counties’ propensity scores, supporting strong overlap. Trimming high- or low-propensity score observations from the sample may be warranted when overlap is weak; for a discussion, see, e.g., Crump_et_al_2009_overlap, Sasaki2022, and Ma2023.
@@
The reliability of these two approaches to estimating conditional DiD designs centers on selecting good models for two different functions: the conditional expectation of untreated outcome changes and the true, unknown propensity score ${p}_{\omega}(X_{i})$. So which approach should one pick? In practice, there are major advantages to combining both in a way that leads to estimators for the $ATT(2)$ that are more robust against model misspecification (SantAnna2020). This is the so-called doubly robust (DR) approach, sometimes referred to as the augmented inverse probability weighting approach; see, e.g., SantAnna2020 and Chang2020.
The key idea of the DR DiD approach is to express the $ATT(2)$ in terms of both $p(X_{i})$ and $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$ in a way that it gives provides some “protection” in case the working models for these functions, $\widehat{\pi}(X_{i})$ and $\widehat{\mu}_{\Delta,G=0}(X_{i})$, are wrong. The resulting DR estimator for the $ATT(2)$ is consistent when either of these nuisance working models is correctly specified. If one is exactly right, it does not matter if the other is wrong. Furthermore, if both working models are only slightly wrong, their errors will multiply, and DR will perform (asymptotically) better than either one alone. [^15] We usually arrive at such estimands by deriving the efficient influence function; see, e.g., SantAnna2020 for a discussion in DiD setups. See also Seaman2018 for an overview. When we adopt nonparametric or machine-learning-based estimators for the nuisance functions, we get a different type of double robustness, “rate double robustness,” where we can trade-off precision between the different working models. For more details, see, Kennedy2017 and Smucler2019. Following the steps in SantAnna2020, we can express the $ATT(2)$ as
@@
with the weights as defined in (4.7). Notice that when we omit the ${m}_{\omega,\Delta,D=0}(X_{i})$ term from (4.10), we are back to the IPW estimand (4.6). When we omit the $w_{\omega,D=0}(D_{i},X_{i})$ term from (4.10), we are back to the regression adjustment estimand (4.2).
Constructing a DR DiD estimator for the $ATT(2)$ based on (4.6) is straightforward: choose a flexible working model for the propensity score and compute its fitted values, $\widehat{\pi}_{\omega}(X_{i})$, a flexible working model for the outcome evolution of untreated units and compute its fitted values for all treated and untreated units, $\widehat{\mu}_{\omega,\Delta,D=0}(X_{i})$, and then use the plug-in principle to estimate the $ATT(2)$ by
@@
We report the $ATT(2014)$ estimates and standard errors using the DR DiD procedure with a linear in covariates working model for ${\mu}_{\omega,\Delta,D=0}(X_{i})$ and (weighted) logistic regression working model that is also linear in covariates for ${p}_{\omega}(X_{i})$ in Table 7 (labeled as “Doubly Robust”). Both unweighted and population-weighted results are fairly similar to IPW DiD estimates. However, an advantage of the DR DiD procedures is that we are arguably less worried about model misspecifications.
Overall, in this section, we discussed different “forward-engineered” estimators for the $ATT(2)$ that respect conditional parallel trends, correctly incorporate covariates, and do not restrict treatment effect heterogeneity. Between RA, IPW, and DR DiD estimators, we recommend practitioners favor the doubly robust one as this procedure adds additional “protection” against modeling assumptions compared to the RA and IPW approaches. When the strong overlap assumption is questionable, though, RA DiD estimators can still work because they extrapolate the outcome model to obtain predicted counterfactual outcome changes even for treated units with covariate values not observed in the comparison group. The credibility of this extrapolation, however, rests on additional assumptions about the accuracy of the working model outside the support of $X_{i}$ in the untreated group. If this extrapolation is not reliable, one can make the DR approach more robust against weak overlap by trimming “extreme” propensity scores and performing a bias correction to ensure that the target parameter (the $ATT(2$) in our case) remains the same. See Ma2023 for details.

### v2

@@ 4.4 DiD estimators with covariates that target the ATT(2)
Fortunately, TWFE is not the only way to bring covariates into DiD estimation. We now discuss alternative ways to use covariates in a DiD analysis that start from the identification result in (4.2):
@@
One way to obtain the second term is to first estimate the inner conditional expectation, $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$. This object is just an (unknown) function that relates average outcome trends for untreated units to their covariates. The most common way to proceed, especially in cases where $X_{i}$ contains many variables or continuous ones, is to specify a working model, ${\mu}_{\omega,\Delta,D=0}(X_{i})$, with parameters that are simple to estimate. A natural and empirically friendly choice is a linear model, ${\mu}_{\omega,\Delta,D=0}(X_{i})=X_{i}^{\prime}\beta_{D=0}$, whose parameters come from a (weighted) regression of $\Delta Y_{i,t=2}$ on $X_{i}$ in the sample of untreated units. The results of this regression describe untreated outcome trends as a function of $X_{i}$. The fitted model then generates predicted values, $\widehat{\mu}_{\omega,\Delta,D=0}(X_{i})=X_{i}^{\prime}\widehat{\beta}_{D=0}$, for all units in the sample, including treated units. With these fitted values we can estimate $\mathop{}\!\mathbb{E}_{\omega}\Big{[}\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]\Big{|}D_{i}=1\Big{]}$ using the plug-in principle; that is, by replacing $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$ with its fitted value $\widehat{\mu}_{\omega,\Delta,D=0}(X_{i})$, and replacing population (weighted) expectations with their sample analogs: $\frac{\sum_{i=1}^{n}D_{i}\leavevmode\nobreak\ \omega_{i}\leavevmode\nobreak\ \widehat{\mu}_{\omega,\Delta,D=0}(X_{i})}{\sum_{i=1}^{n}D_{i}\leavevmode\nobreak\ \omega_{i}}$.
Putting the pieces together gives the following estimator for $ATT(2)$:
@@
This strategy is often referred to as the regression-adjustment (RA) or outcome regression approach to DiD; see, for example, Heckman1997.
We apply this strategy in our application using only the baseline covariates from Table 4, and report the parameters $\widehat{\beta}_{D=0}$ of our working model in columns (1) and (3) of Table 6. In practice, mortality changes in non-expansion states are only weakly related to our baseline covariates. Multiplying the weighted coefficients in Table 6 times the weighted treatment group 2013 means in Table 4 gives a predicted change in untreated mortality rates for the average treated county of 7.2 deaths per 100,000, the second term in equation (4.5), compared to the observed (weighted) change among untreated counties of 6.3 deaths. The observed weighted trend in mortality for expansion counties from Table 2 is 3.7 deaths. Together, these results imply that this approach, based only on assumptions CPT, SO, NA, the choice of baseline covariates, and a linear model for $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$, yields an estimated $ATT(2014)$ of -3.5. This matches the formal RA DiD estimate we report in column 4 of Table 7 (labeled as “Regression”). Column 1, however, gives an unweighted estimate from the same procedure of -1.62.
@@
Unweighted Weighted Regression Propensity Score Regression Propensity Score (1) (2) (3) (4) OLS Logit OLS Logit Constant -20.91 -10.00∗∗∗ -4.62 -8.17∗∗∗ (69.85) (1.35) (44.98) (0.01) % Female 0.04 -0.04∗∗ -0.09 -0.19∗∗∗ (0.82) (0.02) (0.68) (0.00) % White 0.15 0.06∗∗∗ 0.20 0.04∗∗∗ (0.22) (0.00) (0.11) (0.00) % Hispanic -0.08 -0.02∗∗∗ -0.08 -0.02∗∗∗ (0.20) (0.00) (0.08) (0.00) Unemployment Rate 1.14 0.32∗∗∗ 0.88 0.68∗∗∗ (1.56) (0.03) (0.99) (0.00) Poverty Rate 0.21 0.03 -0.13 0.11∗∗∗ (0.98) (0.02) (0.53) (0.00) Median Income 0.09 0.08∗∗∗ -0.05 0.15∗∗∗ (0.52) (0.01) (0.24) (0.00) This table reports the outcome regression propensity score models that enter into the estimator from SantAnna2020 and Callaway2021. The first two columns report the results for unweighted regressions and the second two report results from weighted regression models. The regression model predicts changes in the outcome variable (mortality rates) as a function of 2013 covariate values for the 1,222 counties that do not expand Medicaid in 2014. The propensity score model uses data on all 2,200 counties for 2013 and estimates a logit model of an expansion indicator variable on the 2013 covariate levels. Standard errors (in parentheses) are clustered at the county level.
@@
To compute standard errors, we need to account for the fact that (4.5) is a two-step estimation procedure and take into account the uncertainty associated with estimating the working model $\mu_{\omega,\Delta,D=0}(X_{i})$. This is standard, though, and most statistical software automates this process; see, for example, SantAnna2020 and Callaway2021.
@@
  • This table reports the 2 $\times$ 2 DiD estimate comparing 978 counties in states that expand Medicaid in 2014 to 1,222 counties in states that did not expand Medicaid by 2019, adjusting for 2013 covariate values using the methodologies discussed in SantAnna2020 and Callaway2021. The first column reports results using regression adjustment, the second column uses inverse probability weighting based on a propensity score model using the included covariates, and the third column uses the doubly robust combination of the two approaches. Standard errors (in parentheses) are clustered at the county level.
A linear working model for $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$ is a familiar choice, but not the only one. More flexible, or even fully nonparametric, working models are possible, and the procedure is the same: estimate the model on untreated units, get its fitted values for the covariate values of the treated units, and then estimate the $ATT(2)$ using (4.5). An important consideration when choosing the working models is sample size. Large samples permit more flexible estimators that do not sacrifice too much precision. In smaller samples, a parametric linear model may be more appealing. Ultimately, the reliability of DiD RA estimators for $ATT(2)$ depends on how well $\widehat{\mu}_{\omega,\Delta,D=0}(X_{i})$ approximates $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$. If the working model is misspecified—for example by omitting relevant nonlinear terms—the resulting DiD RA estimator will be biased.
@@
To implement the IPW DiD procedure and construct these “balancing weights,” we need to model $p_{\omega}(X_{i})=P_{\omega}(D_{i}=1|X_{i})$, the (weighted) conditional probability of belonging to the treatment group, known as the “propensity score.” IPW weights are a function of $p_{\omega}(X_{i})$, and for estimating $ATT(2)$, they take a form that forces the underlying weighted distribution of covariates for comparison units to match the distribution for treatment units (Rosenbaum1983). Intuitively, these weights are formed so that if we find some units that were likely to be observed in the treatment groups (based on their covariate values) but ended up in the comparison group, we give these untreated observations “extra” weight.
@@
$ATT(2)$ $=\mathop{}\!\mathbb{E}\left[\Big{(}w_{\omega,D=1}(D_{i})-w_{\omega,D=0}(D_{i},X_{i})\Big{)}\Delta Y_{i,t=2}\right],$ (4.6)
@@
$w_{\omega,D=1}(D)=D\omega\Big{/}{\mathop{}\!\mathbb{E}[\omega D]},\leavevmode\nobreak\ \text{and }\leavevmode\nobreak\ w_{\omega,D=0}(D,X)=\dfrac{\omega(1-D)p_{\omega}(X)}{1-p_{\omega}(X)}\Bigg{/}\mathop{}\!\mathbb{E}\left[\dfrac{\omega(1-D)p_{\omega}(X)}{1-p_{\omega}(X)}\right];$ (4.7)
see, for example, Abadie2005 and SantAnna2020. The structure of the weights in equation (4.7) and the way they enter equation (4.6) highlight several intuitive features of how the IPW estimator works. First, $w_{\omega,D=1}(D)$ is only non-zero for treated units, and $w_{\omega,D=0}(D,X)$ is only non-zero for untreated units, which means the estimator subtracts a particular mean of outcome trends for untreated units from a particular mean of outcome trends for treated units. Second, the $w_{\omega,D=1}(D)$ weights do not involve $X_{i}$ and simply lead simply to a ($\omega$-weighted) mean for the treatment group. Third, the $w_{\omega,D=0}(D,X)$ weights are functions of $X_{i}$ that, as (4.7) shows, give increasingly more weight to untreated units with high propensity scores. This specific weighting function is what allows IPW weights to build a comparison group whose covariate distribution matches the treatment group. Fourth, equation (4.7) clarifies that two types of weights both enter an IPW analysis. We already discussed how the $\omega$ weights shape the parameter of interest, parallel trends assumptions, and estimation. These simply multiply the IPW weights which act to address imbalance (and their product is rescaled to integrate to one within group). Finally, the appeal of (4.6) is that if we have a better sense of how units sort into treatment than of the factors that shape outcome trends, we may be more comfortable modeling $p_{\omega}(X_{i})$ than $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$.
@@
From these fitted values, we can estimate $ATT(2)$ by
@@
We report the $ATT(2014)$ estimates and their standard errors using this IPW DiD procedure in Table 7 (labeled as “IPW”) using both unweighted and population-weighted procedures, where we use a logistic regression that is linear in covariates as our propensity score working model. We use the delta method to account for the estimation uncertainty inherited in this two-step estimation procedure when computing standard errors. The IPW DiD estimates are generally similar to the RA DiD estimates in that we obtain negative estimates that are larger in magnitude when using population weights. However, despite neither being statistically significant, the unweighted IPW estimate is less than half the size of the RA estimate. This is consistent with the broad conclusion from Table 6 that our covariates explain Medicaid expansion better than mortality trends.
IPW estimators for $ATT(2)$ tend to be noisy when fitted propensity score estimates are too close to 1 among untreated units, a condition related to the Assumption SO. The reason for this is simple: when $\widehat{\pi}_{\omega}(X_{i})$ is close to one among units with $D_{i}=0$, the (estimated) inverse probability weights $\widehat{w}_{\omega,D=0}(D,X)$ become more volatile, as one is essentially dividing by zero. A good practice when using IPW estimators is to check the plausibility of strong overlap using estimated propensity scores. Figure 1 plots the propensity scores that come from the logit models in Table 6 for the two groups of counties. Few untreated units have very high estimated propensity scores, so extreme weighting is not a significant concern. [^14] By default, the did and DRDID R packages from Callaway2021 and Santanna2018 trim untreated units with propensity scores in excess of 0.995. In addition, propensity scores of non-expansion counties seem to lie within the support of the expansion counties’ propensity scores, supporting strong overlap. Trimming high- or low-propensity score observations from the sample may be warranted when overlap is weak; for a discussion, see, for example, Crump_et_al_2009_overlap, Sasaki2022, and Ma2023.
@@
The reliability of these two approaches to estimating conditional DiD designs centers on selecting good models for two different functions: the conditional expectation of untreated outcome changes and the true, unknown propensity score ${p}_{\omega}(X_{i})$. So which approach should one pick? In practice, there are major advantages to combining both in a way that leads to estimators for the $ATT(2)$ that are more robust against model misspecification (SantAnna2020). This is the so-called doubly robust (DR) approach, sometimes referred to as the augmented inverse probability weighting approach; see, for instance, SantAnna2020 and Chang2020.
The key idea of the DR DiD approach is to express the $ATT(2)$ in terms of both $p(X_{i})$ and $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$ in a way that it gives provides some “protection” in case the working models for these functions, $\widehat{\pi}(X_{i})$ and $\widehat{\mu}_{\Delta,G=0}(X_{i})$, are wrong. The resulting DR estimator for the $ATT(2)$ is consistent when either of these nuisance working models is correctly specified. If one is exactly right, it does not matter if the other is wrong. Furthermore, if both working models are only slightly wrong, their errors will multiply, and DR will perform (asymptotically) better than either one alone. [^15] We usually arrive at such estimands by deriving the efficient influence function; see, for example, SantAnna2020 for a discussion in DiD setups. See also Seaman2018 for an overview. When we adopt nonparametric or machine-learning-based estimators for the nuisance functions, we get a different type of double robustness, “rate double robustness,” where we can trade-off precision between the different working models. For more details, see, Kennedy2017 and Smucler2019. Following the steps in SantAnna2020, we can express the $ATT(2)$ as
@@
with the weights as defined in (4.7). Notice that when we omit the $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$ term from (4.10), we are back to the IPW estimand (4.6). When we omit the $w_{\omega,D=0}(D_{i},X_{i})$ term from (4.10), we are back to the regression adjustment estimand (4.2).
Constructing a DR DiD estimator for the $ATT(2)$ based on (4.6) is straightforward. Choose a flexible working model for the propensity score and compute its fitted values, $\widehat{\pi}_{\omega}(X_{i})$. Then, a flexible working model for the outcome evolution of untreated units and compute its fitted values for all treated and untreated units, $\widehat{\mu}_{\omega,\Delta,D=0}(X_{i})$, and then use the plug-in principle to estimate the $ATT(2)$ by
@@
We report the $ATT(2014)$ estimates and standard errors using the DR DiD procedure with a linear in covariates working model for ${\mu}_{\omega,\Delta,D=0}(X_{i})$ and (weighted) logistic regression working model that is also linear in covariates for ${p}_{\omega}(X_{i})$ in Table 7 (labeled as “Doubly Robust”). Population-weighted results are fairly similar across our methods, but the unweighted DR DiD estimates are about halfway between the RA and IPW results. However, an advantage of the DR DiD procedures is that model misspecifications are arguably less of a concern.
Overall, in this section, we discussed different “forward-engineered” estimators for the $ATT(2)$ that respect conditional parallel trends, correctly incorporate covariates, and do not restrict treatment effect heterogeneity. Between RA, IPW, and DR DiD estimators, we recommend practitioners favor the doubly robust one compared with the RA and IPW appraoches on their own, as this procedure adds additional “protection.” When strong overlap fails, though, RA DiD estimators can still work because they extrapolate the outcome model to obtain predicted counterfactual outcome changes even for treated units with covariate values not observed in the comparison group. The credibility of this extrapolation, however, rests on the accuracy of the working model outside the support of $X_{i}$ in the untreated group. If this extrapolation is not reliable, one can make the DR approach more robust against weak overlap by trimming “extreme” propensity scores and performing a bias correction to ensure that the target parameter (the $ATT(2$) in our case) remains the same. See Ma2023 for details.